OpenStudy (anonymous):

Find the solutions to the system y=x^2+5x+6 y=4x+12

OpenStudy (anonymous):

\[solve~ x^2+5x+6=4x+12~and~find~the~values~of~x \] and then corresponding values of y.

OpenStudy (ranga):

You are basically substituting the y from the second equation into the first, eliminating y. Solve for x. Bring all the terms on the right in surjithayer's post to the left, simplify, and solve the quadratic equation.

OpenStudy (anonymous):

Solving this system means finding the x -coordinates of the 2 intersections of a line and a parabola. There should be 2 points, meaning two values of x. x^2 + 5x + 6 = 4x + 12 x^2 + x - 6 = 0 D = b^2 - 4ac = 1 + 24 = 25 -> VD = 5 x1 = -1/2 + 5/2 = 2, and x2 = -1/2 - 5/2 = -3.

OpenStudy (ranga):

Solving the quadratic equation by factoring will be a lot easier in this case.

OpenStudy (yanasidlinskiy):

Set the equations equal to each other.

OpenStudy (anonymous):

y=x^2+5x+6 y=4x+12 y=y right? so x^2+5x+6=4x+12 x^2+x-7=0 use quadratic formula and you will have your roots/x coords/x intercepts